I am having trouble finding the volume and surface area of a cone. The lateral side is 10 ft and the base is 9ft.. This is a right cone,don't know if it makes a difference.
Answer by fractalier(6550)

For a right circular cone, the volume is
V = (1/3)(pi)(r^2)h and the surface area (in terms of the lateral side s) is
A = (pi)r^2 + (pi)rs
You haven't specified whether the 9 ft is the radius or diameter, but given these formulas and the one relating height to radius and lateral side, s^2 = r^2 + h^2, you can plug everything in and go...